# Semi-Supervised Unconstrained Head Pose Estimation in the Wild

### v3

@@ 4.6 Other Optional Setups
  Method Unlabeled Dataset $\|\mathbf{I}\!-\!\mathbf{R}_{1}\mathbf{R}_{2}^{T}\|_{F}\downarrow$ Angle error (degree)$\downarrow$
@@
  Method Backbone 2% 5% 10% 20% All
  Sup. TinyViT-22M 4.116 3.734 3.677 3.512 3.455
@@
As discussed in the previous content, in addition to COCOHead [32], there are also other alternative unlabeled datasets such as WiderFace [33], CrowdHuman [34] and OpenImageV6 [31] that contain many wild heads. A natural doubt is, will applying these similar substances yield better results than using COCOHead. To answer this question, we followed the steps of generating COCOHead, processed original WiderFace and CrowdHuman datasets, and obtained the corresponding unlabeled head sets. Specifically, we removed samples with head size smaller than 25, 30 and 30 pixels in WiderFace, CrowdHuman and OpenImageV6. Then, we got about 62K, 163K and 165K heads, respectively. Then, we followed Setting3 to implement the similar DAD-WiderFace, DAD-CrowdHuman and and DAD-OpenImage experiments. Considering that both CrowdHuman and OpenImageV6 have about 2×2$\times$2 × samples than COCOHead ($\sim$74K) or WiderFace, we adjusted their iterations in Phase2 from 100K into 200K. Without loss of generality, all compared models used the ResNet50 [85] as their network backbones.
@@ 5 Conclusion and Discussion
  • Combination with vision-language models. Recently, large vision-language models (VLMs) have demonstrated strong and generalizable visual grounding capabilities by utilizing text descriptions as in large language models (LLMs). Some studies have begun to explore the use of VLMs to address the data scarcity problem [105] and the lack of robust generalization performance [106] in HPE tasks, which is a promising direction worth digging into.
  • More accurate head pose estimation. Our method still cannot address some hard cases shown in Fig. 14, where most facial features are missing due to severe occlusion or backward orientation. We may address these issues by considering the context of other relevant elements, such as the association with body orientation [6] and interaction with surroundings [107, 53].
  • Researches with head pose-prior. Our method is only for the HPE task. We may explore to boost other head-related fields such as 3D face/head reconstruction [103], human object interaction (HOI) [108] and rough eye-gaze estimation [5, 6] by combining the pre-computed head pose information as a priori knowledge.
  • Superior rotation representations. Apart from the introduced probabilistic rotation modeling matrix Fisher, which is widely proven to be superior to the quaternion or ordinary $3\!\times\!3$ rotation matrix, we can also exploit other more robust and generic representations on the rotation manifold, such as Implicit-PDF [109], Rotation NormFlow [110] and Rotation Laplace [72].

### v4

@@ 4.6 Other Optional Setups
  Method
Unlabeled
Dataset
$\|\mathbf{I}\!-\!\mathbf{R}_{1}\mathbf{R}_{2}^{T}\|_{F}\downarrow$
Angle error
[…]
@@
  Type Method Backbone 2% 5% 10% 20% All
  SL Sup. TinyViT-22M 4.116 3.734 3.677 3.512 3.455
@@
As discussed in the previous content, in addition to COCOHead [32], there are also other alternative unlabeled datasets such as WiderFace [33], CrowdHuman [34] and OpenImageV6 [31] that contain many wild heads. A natural doubt is, will applying these similar substances yield better results than using COCOHead. To answer this question, we followed the steps of generating COCOHead, processed original WiderFace and CrowdHuman datasets, and obtained the corresponding unlabeled head sets. Specifically, we removed samples with head size smaller than 25, 30 and 30 pixels in WiderFace, CrowdHuman and OpenImageV6. Then, we got about 62K, 163K and 165K heads, respectively. Then, we followed Setting3 to implement the similar DAD-WiderFace, DAD-CrowdHuman and and DAD-OpenImage experiments. Considering that both CrowdHuman and OpenImageV6 have about $2\times$ samples than COCOHead ($\sim$74K) or WiderFace, we adjusted their iterations in Phase2 from 100K into 200K. Without loss of generality, all compared models used the ResNet50 [87] as their network backbones.
@@ 5 Conclusion and Discussion
  • Combination with vision-language models. Recently, large vision-language models (VLMs) and multimodal LLMs (mLLMs) have demonstrated strong generalizable visual reasoning abilities by aligning image and text inputs. Beyond using text descriptions to mitigate data scarcity [108] or to improve robustness in HPE [109], a more concrete strategy is to cast unconstrained HPE into a visual question answering (VQA) task. For example, both the cropped head image and its reconstructed 3D mesh can be provided to an mLLM, together with a query about the plausibility of the estimated pose. If the mLLM detects a severe error, the system can re-estimate after small perturbations or reject the prediction. Our preliminary test with Gemini 2.5 Flash shows that it can reliably flag most obvious errors, suggesting a promising future direction.
  • More accurate head pose estimation. Our method still cannot address some hard cases shown in Fig. 14, where most facial features are missing due to severe occlusion or backward orientation. A potential remedy is to leverage contextual cues beyond the cropped head. For instance, the orientation of the upper body [6, 110] can provide a strong prior: if the torso clearly faces away from the camera, the head pose should be consistent with a back-of-head orientation. Similarly, surrounding information such as hair-dominated regions with absent facial features [111, 53] can help detect backward heads where the yaw absolute angle must exceed $90^{\circ}$. Integrating such priors may effectively reduce catastrophic errors that cannot be solved by single-frame appearance cues alone.
  • Video-level HPE extension. While our SemiUHPE current focuses on single-frame, applying it to video sequences may lead to temporal jitter or abrupt errors, especially for backward-facing heads. A promising direction is to incorporate temporal consistency constraints, ensuring smooth pose transitions across consecutive frames. Another possibility is to exploit short-term temporal windows, where neighboring frames provide complementary cues to disambiguate extreme poses. Such extensions could further enhance the robustness of SemiUHPE in real-world video applications.
  • Video-based SemiUHPE. Another promising extension is to move from single-frame SemiUHPE to video-based semi-supervised head pose estimation. By redesigning the backbone to process short frame sequences, the model could implicitly capture temporal head-motion cues and output smooth 3D pose trajectories. A teacher–student framework built on annotated video datasets (e.g., BIWI [21], CMU Panoptic [112]) together with large-scale unlabeled online videos would provide a natural baseline. This direction also parallels recent advances in sequential visuomotor imitation learning, where temporal visual observations are mapped to continuous 6DoF trajectories, suggesting that frameworks such as ACT [113] or Diffusion Policy [114] could be adapted for head pose estimation.
  • Impact on downstream tasks. Our SemiUHPE also has implications for broader head-related tasks. For 3D head generation, it can help diagnose and re-balance pose distributions in training datasets, mitigating bias caused by underrepresented large poses. For talking-head generation, robust estimation of extreme angles may reduce jittering artifacts in synthesized videos. For face-swapping, accurate pose priors can guide the alignment of wrapped facial textures, alleviating failures under large yaw angles. These extensions highlight the potential of SemiUHPE as a general building block for robust head-driven applications.